# Binding energies, charge radii, spins and moments: odd-odd Ag isotopes and discovery of a new isomer

### v1

@@ II Experimental methods
In the laser spectroscopy line, the ions were first neutralized using a charge exchange cell filled with potassium. This cell was heated to about 130 degrees Celsius using a NiCr heating wire wrapped around the center, while the ends were kept at 70 degrees Celsius using a liquid cooling loop filled with GALDEN heat transfer liquid. By changing an acceleration potential applied to the cell, the atoms were Doppler-shifted into resonance with the counter-propagating continuous-wave 328 nm laser beam. The laser wavelength was changed in order to keep the scanning voltage between 0 and 2 kV. Measurements were spread over three different experimental periods. For the first experimental run, the laser light was produced using an intra-cavity frequency-doubled Spectra Physics 380 dye laser which was pumped by a 5 W Verdi 532 nm laser. The other measurements used a Matisse Dye laser pumped by a 10 W Millennia 532 nm laser. Instead of intra-cavity doubling, an external WaveTrain frequency-doubler was used. By recording the number of photon counts observed by the photo-multipliers as a function of the wavelength in the rest frame of the atoms, the hyperfine structures of the Ag isotopes could be recorded. To reduce the background from scattered laser light, the recorded number of photons was time-gated such that only the photons measured when the atom bunch was in front of the photo-multiplier were used in the hyperfine spectra. Regular reference measurements were performed on 109Ag by injecting stable beams produced by a dedicated offline spark-discharge ion source Vilén et al. (2020) into the cooler-buncher. Repeated measurements of the same isotopes were performed across all three experiments as an additional systematic check; all three datasets were found to be in mutual agreement. The hyperfine spectra were fitted with the SATLAS analysis suite Gins et al. (2018).
@@
In this work, the systematic uncertainties on the nuclear observables originate from the reference moments and hyperfine - or -constants for the dipole and quadrupole moments, and the field- and mass-shift factor and nuclear masses for the changes in mean-squared charge radii. While the hyperfine anomaly in silver can go up to a few percent Persson (2013), we decided to neglect this contribution as it is not known for the studied isotopes and requires further development. The dipole moment is calculated with a reference = MHz Dahmen and Penselin (1967), = Sahm and Schwenk (1974); Mertzimekis et al. (2016) assuming a zero hyperfine anomaly, and the quadrupole moment with = MHz Fischer et al. (1975) and = b Berkes et al. (1984). The changes in mean-squared charge radii are extracted with newly-calculated atomic field- and mass-shift factors, F = MHz/fm2 and M = GHz u Ohayon et al. (2024), different from the ones reported in Ref. Reponen et al. (2021) based on empirical estimates.
@@ Acknowledgements.
Supplementary Material: New isomeric states in odd-odd Ag isotopes: binding energies, spins, and moments.

### v2

@@ II Experimental methods
In the laser spectroscopy line, the ions were first neutralized using a charge exchange cell filled with potassium. This cell was heated to about 130 degrees Celsius using a NiCr heating wire wrapped around the center, while the ends were kept at 70 degrees Celsius using a liquid cooling loop filled with GALDEN heat transfer liquid. By changing an acceleration potential applied to the cell, the atoms were Doppler-shifted into resonance with the counter-propagating continuous-wave 328 nm laser beam to probe the transition. The laser wavelength was changed in order to keep the scanning voltage between 0 and 2 kV. Measurements were spread over three different experimental periods. For the first experimental run, the laser light was produced using an intra-cavity frequency-doubled Spectra Physics 380 dye laser which was pumped by a 5 W Verdi 532 nm laser. The other measurements used a Matisse Dye laser pumped by a 10 W Millennia 532 nm laser. Instead of intra-cavity doubling, an external WaveTrain frequency-doubler was used. By recording the number of photon counts observed by the photo-multipliers as a function of the wavelength in the rest frame of the atoms, the hyperfine structures of the Ag isotopes could be recorded. To reduce the background from scattered laser light, the recorded number of photons was time-gated such that only the photons measured when the atom bunch was in front of the photo-multiplier were used in the hyperfine spectra. Regular reference measurements were performed on 109Ag by injecting stable beams produced by a dedicated offline spark-discharge ion source Vilén et al. (2020) into the cooler-buncher. Repeated measurements of the same isotopes were performed across all three experiments as an additional systematic check; all three datasets were found to be in mutual agreement. The hyperfine (HF) spectra were fitted with the SATLAS analysis suite Gins et al. (2018).
@@
In the case that the HF structure of multiple nuclear states was overlapped, the highest peak in the HF spectrum was associated with the highest spin nuclear state, as this is the most produced in this experiment Cannarozzo et al. (2023). Moreover, as the number of peaks in the HF structure is larger than the number of free variables (A, A, B and isotope shift ) in 116,118,120Ag, there is no ambiguity in assigning resonances to the difference nuclear states.
In this work, the statistical uncertainties originate from the uncertainty from the fit and the scan-to-scan variation of each fit parameter. The isotope shifts from the lowest spin states, reported in Ref. Reponen et al. (2021), are reported here again, including the scan-to-scan variation. This resulted in consistent values, but larger uncertainties. The systematic uncertainties on the nuclear observables originate from the reference moments and hyperfine - or -constants for the dipole and quadrupole moments, and the field- and mass-shift factor and nuclear masses for the changes in mean-squared charge radii. While the hyperfine anomaly in silver can go up to a few percent Persson (2013), we decided to neglect this contribution as it is not known for the studied isotopes and requires further development. The dipole moment is calculated with a reference = MHz Dahmen and Penselin (1967), = Sahm and Schwenk (1974); Mertzimekis et al. (2016) assuming a zero hyperfine anomaly, and the quadrupole moment with = MHz Fischer et al. (1975) and = b Berkes et al. (1984). The changes in mean-squared charge radii are extracted with newly-calculated atomic field- and mass-shift factors, F = MHz/fm2 and M = GHz u Ohayon et al. (2024), different from the ones reported in Ref. Reponen et al. (2021) based on empirical estimates.
@@ Acknowledgements.
- Cannarozzo et al. (2023) S. Cannarozzo, S. Pomp, A. Solders, A. Al-Adili, A. Göök, and A. Koning, The European Physical Journal A 59, 295 (2023).
@@
Supplementary Material: Binding energies, charge radii, spins and moments: odd-odd Ag isotopes and discovery of a new isomer